40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1889     Entry time: Wed Aug 12 02:00:32 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: report 

Spent a lot of time aligning tonight.  The BS is not staying put--sometimes after a lock loss it gets badly mis-aligned. 

DD handoff is working, after putting beam on REFL diodes and running senseDRM script.

ELOG V3.1.3-